OpenStudy (anonymous):

A population of normally distributed account balances have mean balance $150 and standard deviation of $35. What is the probability that the total balance for a random sample of 40 accounts will exceed $6400?

OpenStudy (campbell_st):

find the average balance for the 40... 6400/40 = $160 this means that the average balance of the 40 selected exceeds $160

OpenStudy (anonymous):

X~N(150,35) X= Total balance / 40 X= 6400/40 X = 160 P(X>160) = ? =P(Z>(1600-150)/35) =P(Z>0.285) = 1- P(z<0.285) =1 - 0.6122 = 0.3878

OpenStudy (campbell_st):

using a z score \[z =(x - \mu)/\sigma\] \[z = (160 - 150)/35\] z = 0.2857

OpenStudy (anonymous):

Thanks! hmm...actually, I have also calculated both the different answers by @DBhatta and @campbell_st. And that's why I am not sure what the correct answer should be. hahaha...

OpenStudy (campbell_st):

I would have thought that if the balance is to exceed then its the probability should be between 50% and 68%... as the average is greater than the mean... and the z score is 0.2857

OpenStudy (anonymous):

I'm pretty sure with my answer and @campbell_st also has the same answer. He has just stopped after calculating z and I've calculated until the end.

OpenStudy (anonymous):

@campbell_st actually it should be below 50%